ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 181, 2011

Alesker, S. A Fourier-type transform on translation-invariant valuations on convex sets

189

Argyros, S. A., I. Deliyanni and A. G. Tolias Hereditarily Indecomposable Banach algebras of diagonal operators

65

Barreira, L. and G. Iommi Multifractal analysis and phase transitions for hyperbolic and parabolic horseshoes

347

Benjamini, I., O. Schramm and S. Sodin Poisson asymptotics for random projections of points on a high-dimensional sphere

381

Burciu, S. Clifford theory for cocentral extensions  111

Conlon, D., J. Fox and B. Sudakov Large almost monochromatic subsets in hypergraphs

423

Deliyanni, I. See S. A. Argyros

Deng, S. On the classification of weakly symmetric Finsler spaces  29

Fox, J. See D. Conlon Ginzburg, D. On the representations of q unipotent groups over a

finite field

387

Giné, J., M. Grau and J. Llibre On the polynomial limit cycles of polynomial differential equations

461

González-Dávila, J. C. and F. Martín Cabrera Harmonic almost contact structures via the intrinsic torsion

145

González-Sánchez, J. and T. S. Weigel Finite p-central groups of height k

125

Grau, M. See J. Giné Iommi, G. See L. Barreira Juschenko, K. and S. Popovych Algebraic reformulation of Connes

embedding problem and the free group algebra

305

Keller, T. M. Finite groups have even more conjugacy classes  433

Llibre, J. See J. Giné Lucchini, A. Profinite groups with nonabelian crowns of bounded

rank and their probabilistic zeta function

53

Martín Cabrera, F. See J. C. González-Dávila Mihăilescu, M. and V. Rădulescu Sublinear eigenvalue problems

associated to the Laplace operator revisited 317 Mityagin, B. An anisotropic integral operator in high temperature

superconductivity

1

Popovych, S. See K. Juschenko Rădulescu, V. See M. Mihăilescu Schramm, O. See I. Benjamini

Shelah, S. Polish Algebras, shy from freedom  477

Sodin, S. See I. Benjamini Solomon, Y. Substitution tilings and separated nets with similarities

to the integer lattice

445

Sudakov, B. See D. Conlon Tolias, A. G. See S. A. Argyros Törnquist, A. Localized cohomology and some applications of Popa’s

cocycle superrigidity theorem

327

Weigel, T. S. See J. González-Sánchez Zweimüller, R. Measure preserving extensions and minimal wander-

ing rates 295